An influenza vaccine is produced by two different companies. It is known that a vaccine produced by company 1 is effective with probability 0.89, while a vaccine produced by company 2 is effective with probability 0.93. We also know that company 1 supplies 40% of the vaccines, while company 2 supplies 60% of the vaccines ordered by the government.

(i) What is the probability that a vaccine is effective, given that it was produced by company 2?

(ii) What is the probability that a randomly chosen vaccine from the government’s order is not effective?

(iii) What is the probability that given a vaccine is not effective that it was produced by company 1?


1
Expert's answer
2021-12-13T14:35:44-0500

Define the following events,

Let "C_1" and "C_2" be the events that the influenza vaccine is manufactured by company 1 and company 2 respectively. Also, let "E" be the event that the influenza vaccine produced is effective and "E'" is the event that the influenza vaccine produced is not effective.

The following probabilities are given,

"p(E|C_1)=0.89,\space p(E|C_2)=0.93,\space p(C_1)=0.40,\space p(C_2)=0.60"


"i)"

The probability that a vaccine is effective, given that it was produced by company 2 is given as,

"p(E|C_2)=0.93".


"ii)"

We determine the probability that a randomly selected vaccine is effective. To do so, we shall apply the law of total probability as follows,

"p(E)=p(E|C_1)*p(C_1)+p(E|C_2)*p(C_2)=0.89*0.40+0.93*0.60=0.356+0.558=0.914"

The probability that a randomly selected vaccine is not effective is given as,

"p(E')=1-p(E)=1-0.914=0.086"

Therefore, the probability that a randomly selected vaccine is not effective is 0.086.


"iii)"

Here, we determine the conditional probability, "p(C_1|E')" defined as,

"p(C_1|E')={p(C_1\cap E')\over p(E')}"

"P(C_1\cap E')=0.4\times0.11=0.044\space and \space p(E')=0.086"

Thus, "p(C_1|E')={0.044\over 0.086}=0.5116(4dp)"

Therefore, the probability that a randomly selected vaccine is produced by company 1 given it is not effective is 0.5116.